OpenStudy (anonymous):

an equation with the roots, 0, -4, -3

OpenStudy (jdoe0001):

\(\bf x=0, \quad x=-4,\quad x=-3\\ \quad \\ \quad \\ x=0\implies x-0=0\implies {\color{blue}{ (x)}}=0\\ \quad \\ x=-4\implies x+4=0\implies {\color{blue}{ (x+4)}}=0\\ \quad \\ x=-3\implies x+3=0\implies {\color{blue}{ (x+3)}}=0\\ \quad \\ {\color{blue}{ (x)(x+4)(x+3)}}=0\implies (x)(x+4)(x+3)=y\\ \quad \\ \implies (x)(x+4)(x+3)=\textit{original polynomial}\)